The volume of an ideal gas varies directly with the temperature T and inversely with the Pressure P. A cylinder contains 15 liters of oxygen at a temperature of 240 K and a pressure of 80 atmospheres. If a piston is lowered into the cylinder decreasing the volume occupied by the gas to 10 liters and the temperature remains constant, what is the pressure?
Answer by venugopalramana(3286)

The volume - V of an ideal gas varies directly with the temperature T and inversely with the Pressure P.
V=K*T/P...WHERE K IS A CONSTANT
A cylinder contains 15 liters of oxygen at a temperature of 240 K and a pressure of 80 atmospheres.
15=K*240/80=3K
K=15/3=5 LITRE.ATM./DEG.K
V=5T/P
If a piston is lowered into the cylinder decreasing the volume occupied by the gas to 10 liters and the temperature remains constant, what is the pressure?
10=5*240/P
P=5*240/10=120 ATM.